# Effectively Controlling Reasoning Models through Thinking Intervention

### v1

@@ Abstract
Reasoning-enhanced large language models (LLMs) explicitly generate intermediate reasoning steps prior to generating final answers, helping the model excel in complex problem-solving. In this paper, we demonstrate that this emerging generation framework offers a unique opportunity for more fine-grained control over model behavior. We propose Thinking Intervention, a novel paradigm designed to explicitly guide the internal reasoning processes of LLMs by strategically inserting or revising specific thinking tokens. We conduct comprehensive evaluations across multiple tasks, including instruction following on IFEval, instruction hierarchy on SEP, and safety alignment on XSTest and SORRY-Bench. Our results demonstrate that Thinking Intervention significantly outperforms baseline prompting approaches, achieving up to 6.7% accuracy gains in instruction-following scenarios, 15.4% improvements in reasoning about instruction hierarchies, and a 40.0% increase in refusal rates for unsafe prompts using open-source DeepSeek R1 models. Overall, our work opens a promising new research avenue for controlling reasoning LLMs.
@@ 1 Introduction
Reasoning-enhanced models, such as OpenAI’s o1 [25], DeepSeek R1 [20], and Google’s Flash Thinking [14], mark a significant advancement in the field of large language models (LLMs). By explicitly incorporating a reasoning process prior to generating responses, these models demonstrate enhanced performance, especially when addressing complex tasks, such as mathematical problem-solving [32], programming [27], and logical reasoning [43].
@@
Despite these developments, the control of reasoning models still primarily relies on traditional input-level operations, such as prompt engineering [41], which modifies the instructions provided to the LLM. In this paper, we demonstrate that the explicit thinking step introduced in reasoning models provides enhanced transparency into the model’s cognitive processes, creating new opportunities for direct and precise interventions within these reasoning stages.
@@
We introduce Thinking Intervention, a novel paradigm designed to explicitly control the internal reasoning processes of models. Rather than allowing the model to generate entire reasoning chains on its own, Thinking Intervention specifies token sequences (e.g., detailed guidance) to be inserted or replaced within the ongoing reasoning process. Such targeted interventions enable fine-grained and transparent control over the reasoning trajectory, aligning the model’s behavior with specific task objectives. Thinking Intervention has the following properties: (1) it does not require any form of model training and can be deployed in real-world settings with minimal engineering effort; (2) it is compatible with existing model control techniques such as prompt engineering and even more sophisticated fine-tuning approaches; (3) it enables fine-grained and flexible control over the reasoning process by adaptively inserting or revising intermediate reasoning steps based on context-specific and task-specific needs.
@@
Demonstration. To further illustrate, consider a general instruction-following task shown in Figure 1 (a), which asks the model to "list 2 famous moms in JSON format". Vanilla Prompting would simply state, "Please provide the names of 2 famous moms in JSON format," while Prompt Engineering might add a reminder, such as "Ensure the format is JSON with 2 famous moms." Although these input-level prompts instruct the model on what to do, reasoning models may still overlook or misunderstand some constraints, as shown in Figure 1(b). In contrast, Thinking Intervention explicitly guides the reasoning by injecting instructions directly into the model’s thinking process, such as "I should generate 2 famous moms and put them in a JSON format." This explicit intervention helps reduce the likelihood of the model missing constraints during the reasoning process, thereby significantly improving performance.
@@
Case studies. We demonstrate the effectiveness of Thinking Intervention across diverse tasks, including instruction following on IFEval, instruction hierarchy on SEP, and safety alignment on XSTest and SORRY-Bench benchmarks. For instruction following (Section 3), we illustrate how Thinking Intervention can guide the model to follow instructions more accurately and effectively. For instruction hierarchy (Section 4), we show that Thinking Intervention helps the model reason about hierarchical instructions and appropriately prioritize main tasks over low-priority ones. As shown in Figure 1(b), Thinking Intervention significantly outperforms both the Vanilla Prompting and Prompt Engineering methods, achieving 6.7% and 1.9% improvements in instruction following tasks and 15.4% and 4.0% improvements in instruction hierarchy tasks, respectively.
Lastly, in the safety alignment task (Section 5), we first reveal that open-source reasoning models (e.g., the DeepSeek R1 model series [20]) tend to over-comply with user instructions, even when those instructions pose safety risks. This behavior underscores the urgent need for more effective safety steering methods. We then demonstrate that applying Thinking Intervention enables explicit guidance toward safer reasoning patterns, leading to substantial improvements in safety performance. For example, as shown in Figure 1(b), Thinking Intervention significantly increases the refusal rates for unsafe requests by up to 40.0% and 26.9% on the XSTest and SORRY-Bench, respectively.
@@
Overall, our work demonstrates that the Thinking Intervention, as a novel paradigm, provides a powerful and flexible approach to enhance the capabilities of reasoning models across multiple dimensions. We encourage the community to explore and extend this framework for various applications, as it opens up new possibilities for precise, transparent, and effective control over the reasoning processes of LLMs, ultimately leading to more reliable and aligned AI systems.
2 Thinking Intervention: A Novel Perspective on Controlling Reasoning Models
@@ 2.2 Intervening in the Reasoning Process as a General Paradigm
Traditional approaches111Fine-tuning can also be viewed as a form of model control in certain cases, but we do not consider it here because it is more destructive—it modifies the entire model rather than controlling a fixed one. to improving LLM performance have largely focused on prompt engineering [57, 54, 64, 39], which optimizes the input context to elicit better model response . For reasoning-enhanced LLMs, although crafting the initial prompt remains important, the explicit reasoning stage offers a new, more direct pathway for optimization: intervening within the reasoning process itself.
In this work, we propose a general paradigm termed Thinking Intervention, which directly intervenes within the reasoning process of LLMs, e.g., through revising explicit instructions or guidance at intermediate reasoning steps. Unlike prompt engineering, where the entire input context is optimized before the LLM generates tokens, Thinking Intervention operates in an online, dynamic environment. The reasoning chain is generated token-by-token in real time, requiring the intervention mechanism to make decisions based on the incomplete reasoning chain . The key challenge lies in developing intervention strategies that can quickly evaluate partially completed reasoning paths, intervene at appropriate intermediate steps, and conduct intervention strategies adapted to the diverse reasoning trajectories while not disrupting the model’s natural reasoning flow.
@@ 2.3 Instantiation: Intervention via a Postfix-based Monitor
A simple yet powerful instantiation of the intervention function is based on monitoring the reasoning chain to detect specific trigger strings. Specifically, given a set of trigger strings (which can be single tokens or sequences of multiple tokens), the monitor checks if the most recent tokens (i.e., the postfix of the current reasoning chain) match any string in . If a match is detected, we append an intervention sequence (e.g., "I am a responsible assistant") immediately to the existing reasoning chain. The trigger set can be designed flexibly to capture relevant reasoning stages, domain-specific phrases, or other reasoning markers. Formally, the intervention function with a postfix-based monitor is defined as:
Here, denotes concatenation of the current reasoning chain with the intervention sequence . Various practical intervention strategies can be readily implemented by appropriately selecting trigger strings . Below, we illustrate several simple, easy-to-implement examples, though more sophisticated approaches are also possible:
Intervention at reasoning start. To intervene at the beginning of reasoning, a trigger string indicating reasoning onset (e.g., "<think>" in DeepSeek R1 models) can be included in . Upon detecting this string, the model immediately receives guidance through an intervention sequence containing relevant instructions or hints to direct reasoning from the start.
Intervention at reasoning conclusion. Trigger strings signaling the end of the reasoning phase (e.g., "</think>" in DeepSeek R1 models) can be included in . Detecting these triggers allows reinforcing key points or addressing overlooked considerations before finalizing the output.
Intervention at reasoning transitions. To intervene at intermediate stages of reasoning, trigger strings corresponding to transitional markers or reflective cues (e.g., "wait", "Hmm", "Alternatively") can be specified. Interventions triggered at these junctures prompt the model to reconsider previous steps, correct potential errors, or further elaborate on its reasoning, thereby enhancing the quality and rigor of the resulting reasoning chain.
In our evaluations, we experimented with all three strategies and observed that intervention at the beginning of the reasoning process is the most effective. Therefore, throughout our experiments, we primarily focused on using this approach. We believe this is because the initial reasoning tokens are particularly important and can directly guide the reasoning trajectory toward more productive outcomes.
@@ 2.4 Thinking Intervention enjoys several unique features and advantages
Flexibility in design. The generation of Thinking Intervention sequences can be designed manually by domain experts or automatically synthesized by auxiliary models. For example, users or LLM providers can use another LLM to transform key instructions into intervention sequences. Additionally, Thinking Intervention can be injected or edited multiple times at arbitrary positions throughout the reasoning process, offering a versatile and robust framework for precise control over reasoning paths.
Simplicity and lightweight. Thinking Intervention is straightforward and easy to implement in any reasoning model. It does not require additional fine-tuning or reinforcement learning to adjust model behavior. Moreover, it incurs no significant computational cost, as it only involves adding or editing a few tokens within the reasoning process. If the intervention guides the model into correct reasoning paths at an early stage, it can even reduce the number of reasoning tokens needed to arrive at the final answer, thus saving computation time.
Compatibility with other methods. Thinking Intervention can be combined with other techniques, such as prompt engineering, to further enhance model performance. For example, we can first use prompt engineering to provide the model with some background knowledge and then use Thinking Intervention to guide the model’s reasoning process. This combination allows for a more comprehensive approach to controlling reasoning models.
First-person narrative. One interesting property that distinguishes Thinking Intervention from prompt engineering is its use of the first-person narrative. For instance, while prompt engineering might use "You are a responsible assistant", Thinking Intervention would instead use "I am a responsible assistant", as it is integrated into the reasoning process. This first-person narrative allows the model to treat the instruction as part of its own reasoning rather than as an external command.
3 Case Study: Instruction Following
@@
In this section, we demonstrate how our proposed Thinking Intervention method effectively improves the instruction-following capability of reasoning models.
3.1 Experimental Design
@@
Benchmark. We leverage Instruction-Following Evaluation (IFEval) [69] as a benchmark to measure how well reasoning models follow instructions. The benchmark comprises 500 prompts, each containing one or more of 25 identified types of verifiable instructions. These verifiable instructions function as constraints on the output, such as "do not use any commas" or "no capital letters are allowed" To quantify instruction following capability, we report the accuracy, defined as the proportion of prompts for which the model satisfies all verifiable instructions within the prompt. Additional evaluation details are provided in Appendix B.1.
3.2 Methods
Baseline prompting methods. We evaluate two common baseline techniques: Vanilla Prompting and Reminder Prompting. Specifically, Vanilla Prompting directly feeds the complete prompt into the reasoning model without modification. In contrast, Reminder Prompting augments the original prompt by explicitly reiterating the constraint. To generate the corresponding reminder, we prompt an auxiliary LLM with the instructions from IFEval. The full prompt is provided in Appendix B.1.
Generating Thinking Intervention. To ensure that constrained instructions are correctly followed, we use Thinking Intervention to explicitly highlight the importance of these constraints during the reasoning process. To construct the intervention sequence , we leverage the Reminder Prompting generated by LLMs and make slight edits to it. For instance, given an original instruction such as "without using any commas", the LLM may produce a reminder like "Ensure that you do not use any commas." We then apply light narrative editing to transform this into the intervention sequence, such as "I should not use any commas." This intervention sequence is then inserted at the beginning of the reasoning process throughout the evaluations.
We illustrate a concrete example of how Thinking Intervention can be integrated with Vanilla Prompting and Reminder Prompting in Figure 2.
| (a) R1-Qwen-14B | (b) R1-Qwen-32B | (c) QwQ-32B |
3.3 Results
The evaluation results on the IFEval dataset, summarized in Figure 3, demonstrate the effectiveness of Thinking Intervention across multiple reasoning models (R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B). Additional experimental results related to more evaluation metrics are provided in Appendix B.2.
Thinking Intervention consistently improves instruction-following capability. We observe that Vanilla Prompting with Thinking Intervention yields the most significant gains across all three models, increasing accuracy by 4.99%, 6.65%, and 2.96% for R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B, respectively. We also observe that optimal performance is consistently achieved by incorporating Thinking Intervention. For R1-Qwen-14B, combining Thinking Intervention with Vanilla Prompting yields the best results, while for R1-Qwen-32B and QwQ-32B, pairing Thinking Intervention with Reminder Prompting achieves the highest accuracies of 77.63% and 82.44%, respectively, highlighting the significance of Thinking Intervention.
These consistent improvements across different models demonstrate that Thinking Intervention effectively guides model reasoning by explicitly emphasizing constrained instructions, providing a robust enhancement to existing prompting strategies without requiring additional training.
4 Case Study: Instruction Hierarchy
@@
Next, we explore how Thinking Intervention benefits the instruction hierarchy task [53, 59], which evaluates a model’s ability to appropriately prioritize high-priority instructions while disregarding low-priority ones. This capability is essential for safety-critical applications, where models must adhere to specific guidelines even in the presence of conflicting instructions. We examine how Thinking Intervention enhances the model’s ability to navigate complex scenarios involving competing directives.
4.1 Experimental Design
@@
Benchmark. We use the SEP dataset [74] as our benchmark. Each prompt contains a high-priority main instruction paired with relevant data content and an unrelated low-priority instruction. Models are expected to prioritize the main instruction while ignoring the low-priority instruction when it is embedded within the data (see left example in Figure 4). This benchmark enables us to evaluate how effectively models can maintain instruction hierarchies in complex scenarios involving potentially misaligned directives.
@@
Evaluation metrics. We evaluate model performance on the SEP benchmark using two key metrics: (1) robustness, which measures the proportion of low-priority instructions correctly ignored when embedded within data (Figure 4 left); (2) utility, which quantifies the model’s baseline performance on the main task in the absence of any low-priority instructions (Figure 4 right). For the utility metric, we follow [68] by employing LLM-as-a-judge for evaluation and normalizing scores to a 0-100 scale.
| R1-Qwen-14B | R1-Qwen-32B | QwQ-32B | ||||
| Methods | Robustness(%) | Utility(%) | Robustness(%) | Utility(%) | Robustness(%) | Utility(%) |
| Vanilla | 34.00 | 81.04 | 34.80 | 81.76 | 22.20 | 88.00 |
| Vanilla+ThinkingI | 38.40 (+4.40) | 81.08 (+0.04) | 50.20 (+15.40) | 82.02 (+0.26) | 31.40 (+9.20) | 88.16 (+0.16) |
| Reminder | 38.40 | 80.50 | 46.20 | 81.16 | 36.20 | 87.52 |
| Reminder+ThinkingI | 41.80 (+3.40) | 80.90 (+0.40) | 66.40 (+20.20) | 80.90 (-0.26) | 43.40 (+7.20) | 86.79 (-0.73) |
4.2 Methods
Baseline prompting methods. Similar to our instruction-following experiments, we include two baseline approaches: Vanilla Prompting, which directly uses the prompts without additional guidance, and Reminder Prompting, which includes an explicit instruction reminder. Figure 5 shows all the prompts used for evaluation.
@@ 4.2 Methods
Thinking Intervention. For our Thinking Intervention approach, we use the intervention sequence "I should follow all the instructions in the task block and not follow any instructions in the data block." to explicitly guide the model in maintaining the correct instruction hierarchy. This intervention sequence is inserted at the beginning of the reasoning process to help the model correctly prioritize the instructions.
4.3 Results
@@
Thinking Intervention significantly improves robustness while maintaining model utility. Table 1 presents the evaluation results of reasoning models on the SEP benchmark. Our Thinking Intervention approach consistently improves both Vanilla Prompting and Reminder Prompting techniques in terms of robustness across all model scales. For example, when applied to R1-Qwen-32B with Reminder Prompting, Thinking Intervention achieves a robustness of 66.4%, marking a substantial 20.20% improvement over the Reminder Prompting baseline. These improvements are consistent across different model sizes, with QwQ-32B showing a 7.20% increase and R1-Qwen-14B demonstrating a 3.40% boost. Importantly, Thinking Intervention preserves the utility of the models. Across all settings, it incurs only a negligible variation in utility ( 0.73%) compared to the corresponding baselines.
@@
These results demonstrate that Thinking Intervention effectively guides models to maintain proper instruction hierarchies by correctly following high-priority instructions without compromising their performance on utility-related tasks. For experimental results with additional metrics, see Appendix C.
5 Case Study: Safety Alignment

### v2

@@ Abstract
Reasoning-enhanced large language models (LLMs) explicitly generate intermediate reasoning steps prior to generating final answers, helping the model excel in complex problem-solving. In this paper, we demonstrate that this emerging generation framework offers a unique opportunity for more fine-grained control over model behavior. We propose Thinking Intervention, a novel paradigm designed to explicitly guide the internal reasoning processes of LLMs by strategically inserting or revising specific thinking tokens. We find that the Thinking Intervention paradigm enhances the capabilities of reasoning models across a wide range of tasks, including instruction following on IFEval, instruction hierarchy on SEP, and safety alignment on XSTest and SORRY-Bench. Our results demonstrate that Thinking Intervention significantly outperforms baseline prompting approaches, achieving up to 6.7% accuracy gains in instruction-following scenarios, 15.4% improvements in reasoning about instruction hierarchies, and a 40.0% increase in refusal rates for unsafe prompts using open-source DeepSeek R1 models. Overall, our work opens a promising new research avenue for controlling reasoning LLMs. WARNING: This paper contains red-teaming content that can be offensive.
@@ 1 Introduction
Recent advances in reasoning-enhanced models, including OpenAI’s o1 [27], DeepSeek’s R1 [20], and Google’s Flash Thinking [13], have significantly expanded the capabilities of large language models (LLMs). By explicitly incorporating intermediate reasoning steps before producing answers, these models exhibit substantially improved performance in handling complex tasks such as mathematical problem-solving [36], programming assistance [29], and logical inference [48].
@@
Despite these developments, existing methods for guiding reasoning models still predominantly rely on input-level manipulations, such as prompt engineering [46], which modifies the instructions provided to the LLM. In this work, we demonstrate that the explicit thinking steps introduced in reasoning LLMs not only provide enhanced transparency into the model’s cognitive processes but also create new opportunities for direct and precise interventions within these reasoning stages.
@@
Motivated by this insight, we introduce Thinking Intervention, a novel paradigm that explicitly controls the internal reasoning processes of models. Rather than allowing the model to generate entire reasoning chains on its own, Thinking Intervention specifies token sequences (e.g., detailed instructions, clarifications, or constraints) to be inserted or replaced within the ongoing reasoning process. Such targeted interventions enable fine-grained and transparent control over the reasoning trajectory, closely aligning the model’s behavior with required task objectives.
@@
Demonstration. To further illustrate, consider a general instruction-following task shown in Figure 1(a), which asks the model to "list 5 famous moms in JSON format." A standard Vanilla Prompting would just state the instruction and Prompt Engineering might add a reminder like "Ensure the format is JSON." Nonetheless, reasoning models may still overlook essential instructions or constraints (Figure 1(b)). In contrast, Thinking Intervention explicitly guides the reasoning process by injecting instructions directly into the model’s internal thought process, e.g., "I should generate 5 famous moms and put them in a JSON format." Such precise intervention reduces the likelihood of the model missing constraints during reasoning, thereby significantly improving performance.
Notably, Thinking Intervention presents several key advantages: First, it enables fine-grained and flexible control over the reasoning process by adaptively inserting or revising intermediate reasoning steps based on task-specific needs; second, it does not require any form of model training and can be deployed in real-world settings with minimal engineering effort; and third, it seamlessly integrates and complements existing techniques, including prompt engineering and advanced fine-tuning methods.
@@
Empirical findings. We demonstrate the effectiveness of Thinking Intervention across diverse tasks, including instruction following on IFEval [77], instruction hierarchy on SEP [81], and safety alignment on XSTest [45] and SORRY-Bench [68]. For instruction following tasks (§3), we show that applying Thinking Intervention enables the model to more effectively and accurately follow task instructions, leading to notable improvements of 6.7% and 1.9% over baseline Vanilla Prompting and Prompt Engineering methods respectively (Figure 1(b)). Similarly, in instruction hierarchy scenarios (§4), Thinking Intervention guides the model to reason about hierarchical instructions and appropriately prioritize main tasks over low-priority ones, thereby boosting robustness up to 15.4% compared to baselines (Figure 1(b)). Lastly, in safety alignment tasks (§5), we first show that open-source reasoning models (e.g., DeepSeek R1 [20]) often over-comply with unsafe instructions, highlighting an urgent need for better safety control methods. We then demonstrate that Thinking Intervention explicitly steers models toward safer reasoning, substantially increasing refusal rates for unsafe requests by up to 40.0% on XSTest and 26.9% on SORRY-Bench (Figure 1(b)).
Finally, in Section 6, we analyze internal model mechanisms (e.g., attention maps) and find that the attention during the reasoning process predominantly focuses internally rather than on external input tokens, explaining why Thinking Intervention achieves greater effectiveness. We also discuss practical design considerations for implementing Thinking Intervention in real-world settings.
@@
Overall, our findings establish Thinking Intervention as a powerful, flexible, and broadly applicable paradigm for enhancing reasoning model across multiple dimensions. We encourage the community to further explore and adopt this framework, which offers precise, transparent, and effective control over LLM reasoning processes, ultimately enabling more reliable and aligned AI systems.
2 Thinking Intervention: A Novel Paradigm to Control Reasoning Models
@@ 2.2 Intervening in the Reasoning Process as a General Paradigm
Traditional approaches111Fine-tuning can also be viewed as a form of model control in certain cases, but we do not consider it here because it is more destructive—it modifies the entire model rather than controlling a fixed one. to improving LLM performance have largely focused on prompt engineering [64, 60, 71, 44], which optimizes the input to elicit better response . For reasoning-enhanced LLMs, although crafting the initial prompt remains important, the explicit reasoning stage offers a new, more direct pathway for optimization: intervening within the reasoning process itself.
In this work, we propose a general paradigm termed Thinking Intervention, which directly intervenes within the reasoning process of LLMs, e.g., through revising explicit instructions or guidance at intermediate reasoning steps. Unlike prompt engineering, where the input context is optimized before generating tokens, Thinking Intervention operates in an online, dynamic environment. The reasoning chain is generated token-by-token in real time, requiring the intervention mechanism to make decisions based on the incomplete reasoning chain . The key challenge lies in developing intervention strategies that can quickly evaluate partial reasoning paths, intervene appropriately, and adapt to current trajectories without disrupting the LLM’s natural reasoning flow.
@@ 2.3 Instantiation: Intervention via a Postfix-based Monitor
A simple yet powerful instantiation of the intervention function is based on monitoring the reasoning chain to detect specific trigger strings. Specifically, given a set of trigger strings (which can be single tokens or sequences of multiple tokens), the monitor checks if the most recent tokens (i.e., the postfix of the current reasoning chain) match any string in . If a match is detected, we append an intervention sequence (e.g., "I am a safe and responsible assistant") immediately to the existing reasoning chain. The trigger set can be designed flexibly to capture relevant reasoning stages, domain-specific phrases, or other reasoning markers. Formally, the intervention function with a postfix-based monitor is defined as:
Here, denotes the concatenation of the current reasoning with intervention sequence . Practical intervention strategies can be efficiently implemented by selecting appropriate triggers . We illustrate three easy-to-implement examples, though more sophisticated approaches are feasible. (1) Intervention at reasoning start: To intervene at the beginning of reasoning, a trigger string indicating reasoning onset (e.g., "<think>" in DeepSeek R1 models) can be included in . Upon detecting this string, the model immediately receives guidance through an intervention sequence containing relevant instructions or hints to direct reasoning from the start. (2) Intervention at reasoning conclusion: Utilize end-of-reasoning trigger strings (e.g., "</think>") within to reinforce critical points or identify overlooked issues before generating the final output. (3) Intervention at reasoning transitions: Incorporate transitional markers (e.g., "wait", "Hmm") into to prompt the model to review previous reasoning steps, correct potential mistakes, or provide further elaborations.
In our evaluations, we experimented with all three strategies and found that intervening at the beginning of the reasoning process was the most effective. Consequently, we primarily adopted this approach throughout our main experiments (§3, §4, §5) and analyzed alternative strategies in the discussion section (§6.2). In addition to the postfix-based intervention, we also implemented an adaptive strategy that leverages an auxiliary LLM to revise reasoning traces, as presented in Section 6.2.
@@ 2.4 Thinking Intervention enjoys several unique features and advantages
Design flexibility. The intervention sequence can be manually designed by domain experts or automatically synthesized using auxiliary models (e.g., LLMs translating high-level task requirements into specific targeted interventions). Interventions may also be instruction-independent (e.g., for safety alignment tasks) or instruction-dependent (e.g., to enhance specific instruction-following), allowing broad applicability. Finally, interventions can be flexibly applied at arbitrary positions throughout the reasoning process, enabling versatile control over the model’s reasoning.
Easy integration. Adopting Thinking Intervention requires minimal engineering effort and, crucially, does not necessitate fine-tuning or modifying the underlying model parameters. Furthermore, the intervention also incurs negligible computational overhead, ensuring lightweight deployment.
Broad compatibility. Thinking Intervention naturally integrates with established approaches such as prompt engineering, forming synergistic combinations to maximize effectiveness. For instance, prompt engineering can supply background information, while Thinking Intervention can explicitly guide intermediate reasoning steps, thereby enhancing overall model performance and consistency.
Effectiveness. Thinking Intervention effectively controls models by directly guiding the reasoning process, ensuring key instructions are actively integrated rather than passively referenced. As we will show later in Section 6, Thinking Intervention elicits stronger model attention on intervention tokens compared to prompt engineering (Figure 7). This leads to consistent improvements across instruction following (§3), instruction hierarchy (§4), and safety alignment tasks (§5).
3 Evaluation on Instruction Following Tasks
@@
In this section, we empirically demonstrate how our proposed Thinking Intervention approach significantly enhances the instruction-following capability of reasoning models.
| (a) R1-Qwen-7B | (b) R1-Qwen-14B | (c) R1-Qwen-32B | (d) QwQ-32B |
@@
Benchmark and models. We leverage Instruction-Following Evaluation (IFEval) [77] to measure how well reasoning models follow instructions. The benchmark comprises 500 prompts, each containing some identified types of verifiable instructions. These verifiable instructions function as constraints on the output, such as "do not use any commas". To quantify the model’s capability, we report the accuracy, defined as the proportion of prompts for which the responses satisfy all verifiable instructions within the prompt. We consider reasoning models distilled from DeepSeek R1 [20], including R1-Qwen-7B, R1-Qwen-14B, and R1-Qwen-32B, and the QwQ-32B model [43]. Further results on R1-Llama models and evaluation details are provided in Appendix B. All models use the same tag, <think> and </think>, to denote the start and end of the reasoning process.
Methods. We compare Thinking Intervention against two baselines: Vanilla Prompting, which directly uses original prompts unchanged, and Reminder Prompting, which augments the input prompts by restating instruction constraints as reminders. These reminders are generated by prompting an auxiliary LLM using original instructions. For Thinking Intervention, we convert reminder statements into first-person narrative interventions (i.e., changing "Ensure the answer..." to "I should ensure the answer..." via a fixed prefix)222In practice, intervention sequences can be directly constructed using auxiliary LLMs.. This intervention sequence is then inserted at the beginning of the reasoning process. Figure 2 (top) illustrates how Thinking Intervention integrates with the Vanilla Prompting and Reminder Prompting.
Thinking Intervention consistently improves instruction-following capability. Figure 2 (bottom) illustrates consistent improvements across multiple reasoning models. Specifically, integrating Thinking Intervention with the baseline Vanilla Prompting yields accuracy gains of 5.74%, 4.99%, 6.65%, and 2.96% for R1-Qwen-7B, R1-Qwen-14B, R1-Qwen-32B, and QwQ-32B, respectively. Notably, the effectiveness of Thinking Intervention is preserved as the model size increases within the R1-Qwen family, with R1-Qwen-32B showing even greater benefits than R1-Qwen-7B and R1-Qwen-14B. Further performance improvement is achieved by combining Thinking Intervention with the Reminder Prompting method: accuracy reaches as high as 62.84% (R1-Qwen-7B), 77.63% (R1-Qwen-32B), and 82.44% (QwQ-32B). Thus, Thinking Intervention not only provides substantial stand-alone gains across model families, but also complements existing prompting methods.
Overall, these findings confirm Thinking Intervention’s capability to precisely guide reasoning models to follow constrained instructions. Crucially, this significant performance boost is achieved without requiring additional model training. Moreover, we explore the broader applicability of Thinking Intervention in Appendix C, showing how explicit intervention during the reasoning process can mitigate model overthinking, thereby highlighting the versatility and utility of the proposed method.
4 Evaluation on Instruction Hierarchy Task
@@
Next, we explore how Thinking Intervention benefits the instruction hierarchy task [59, 66], which evaluates a model’s ability to appropriately prioritize high-priority instructions over low-priority ones. This capability is essential for safety-critical applications, where models must adhere to specific guidelines even in the presence of conflicting instructions. We examine how Thinking Intervention enhances the model’s ability to navigate complex scenarios involving competing directives.
| R1-Qwen-7B | R1-Qwen-14B | R1-Qwen-32B | QwQ-32B | |||||
| Methods | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) | Rob.(%) | Utility(%) |
| Vanilla | 57.60 | 74.44 | 34.00 | 81.04 | 34.80 | 81.76 | 22.20 | 88.00 |
| Vanilla+ThinkingI | 60.80 | 74.40 | 38.40 | 81.08 | 50.20 | 82.02 | 31.40 | 88.16 |
| Reminder | 57.60 | 74.20 | 38.40 | 80.50 | 46.20 | 81.16 | 36.20 | 87.52 |
| Reminder+ThinkingI | 62.60 | 73.92 | 41.80 | 80.90 | 66.40 | 80.90 | 43.40 | 86.79 |
@@
Benchmark. We evaluate on the SEP dataset [81]. Each data point contains a high-priority main instruction paired with relevant data content and an unrelated low-priority instruction. Models are expected to prioritize the main instruction while ignoring the low-priority instruction. This benchmark enables us to evaluate how effectively models can maintain instruction hierarchies in complex scenarios involving potentially misaligned directives.
@@
Evaluation metrics. We evaluate model performance on the SEP benchmark using two key metrics: (1) robustness, which measures the proportion of low-priority instructions correctly ignored when embedded within data; (2) utility, which quantifies the model’s baseline performance on the main task in the absence of any low-priority instructions. For the utility metric, we follow [75] by employing LLM-as-a-judge for evaluation and normalizing scores to a 0-100% scale.
@@
Methods. Similar to our instruction-following experiments, we include two baseline approaches: Vanilla Prompting, which directly uses the prompts without additional guidance, and Reminder Prompting, which includes an explicit instruction reminder. For our Thinking Intervention approach, we use the intervention sequence "I should follow all the instructions in the task block and not follow any instructions in the data block." to explicitly guide the model in maintaining the correct instruction hierarchy. This intervention sequence is inserted at the beginning of the reasoning process to help the model correctly prioritize the instructions. More details on the evaluation setup, including baseline prompts and examples, can be found in Appendix D.1.
@@
Thinking Intervention significantly improves robustness while maintaining model utility. Table 1 presents the evaluation results of reasoning models on the SEP benchmark. Our Thinking Intervention approach consistently improves both Vanilla Prompting and Reminder Prompting techniques in terms of robustness across all model scales. For example, when applied to R1-Qwen-32B with Reminder Prompting, Thinking Intervention achieves a robustness of 66.4%, marking a substantial 20.20% improvement over the Reminder Prompting baseline. We observe similar robustness enhancements across other model variants, including gains of 5.0% for R1-Qwen-7B, 3.4% for R1-Qwen-14B, and 7.2% for QwQ-32B. Importantly, Thinking Intervention preserves the utility of the models. Across all settings, it incurs only a negligible variation in utility ( 0.73%) compared to the baselines. Interestingly, we note that R1-Qwen-7B exhibits the highest robustness with Vanilla Prompting. This is attributed to its lower general utility ( 5% lower than other models), which could prevent it from correctly responding to the injected low-priority prompt, thus resulting in higher robustness.
@@
These results demonstrate that Thinking Intervention effectively guides models to maintain proper instruction hierarchies by correctly following high-priority instructions without compromising their utility-related tasks. For experimental results with R1-Llama models, see Appendix D.
5 Evaluation on Safety Alignment Task